production possibilities

The following table describes the production possibilities of two state in the country of Malaysia:

 


Wheat per Worker per Hour

Corn per Worker per Hour

Kedah

6

6

Perlis

2

4

 

(a)   Without trade, what is the price of wheat (in terms of corn) in Kedah? What is the price in Perlis?


(b)   Which state has an absolute advantage in the production of wheat and corn? Which state has a comparative advantage in the production of wheat and corn? Explain your answer.
Answer #1

Opportunity cost of producing 1 pair of red socks (Boston) = 3/3 = 1 pair of white socks

Opportunity cost of producing 1 pair of red socks (Chicago) = 1/2 = 0.5 pair of white socks

a)

Price of 1 pair of white socks in Boston = 1 pair of red socks

Price of 1 pair of white socks in Chicago = 1/0.5 = 2 pair of red socks

b)

Boston has an absolute advantage in production of both socks. (more pairs can be produced per worker per hour)

Boston has a comparative advantage in production of white socks while Chicago has a comparative advantage in production of red socks. (lower opportunity cost)


answered by: ANURANJAN SARSAM